Johnny gets to work in 1/2 hour when he drives his car. Riding his bike (by the same route) it takes him 1 hr. His average driving speed is 4.5 mph greater than his average speed on his bike. What is his average speed by car, and by bike? How far does he travel to work?
Answer by ankor@dixie-net.com(22740)

Johnny gets to work in 1/2 hour when he drives his car.
Riding his bike (by the same route) it takes him 1 hr.
His average driving speed is 4.5 mph greater than his average speed on his bike.
What is his average speed by car, and by bike? How far does he travel to work?
:
Let s = driving speed by car
then
(s-4.5) = riding speed on bike
:
Write a distance equation: Dist = time * speed
By bike = by car
1(s-4.5) = .5s
s - 4.5 = .5s
s - .5s = +4.5
.5s = 4.5
s = 9 mph is the car
then
9 - 4.5 = 4.5 mph by bike